Find the short leg of a 30-60-90 triangle,if the hypotenuse measures 7 feet.
Answer by CharlesG2(834)

Find the short leg of a 30-60-90 triangle,if the hypotenuse measures 7 feet.
sohcahtoa --> sin = opp/hyp, cos = adj/hyp, tan = opp/adj
pythagorean theorem --> a^2 + b^2 = c^2, where a and b are legs, and c is the hypotenuse
sin 30 = opp/7
0.5 = opp/7
7 * 0.5 = opp
3.5 = opp --> short leg
cos 30 = adj/7
sqrt(3)/2 = adj/7
7sqrt(3)/2 = adj
6.06218 approx. = adj --> long leg
check:
(7/2)^2 + (7sqrt(3)/2)^2 = c^2
49/4 + (49*3)/4 = c^2
(49 + 147)/4 = c^2
196/4 = c^2
49 = c^2
7 = c